OpenStudy (anonymous):

Need help on (x - 4)(x + 5) = 7?

OpenStudy (anonymous):

solve both terms for x. So, x-4=7 x+5=7 solve both for x.

OpenStudy (anonymous):

Expand the brackets. How? Distribute EACH term. (x - 4)(x + 5) = 7 x * x = x^2 x * 5 = 5x -4 * x = -4x -4 * 5 = -20 So now you have: x^2 + 5x - 4x - 20 = 7 Combine like terms. Solve for x. x^2 + 5x - 4x - 20 = 7 x^2 + x - 20 = 7 Subtract 7. x^2 + 5x - 13 = 0 ^That is a quadratic equation (ax^2 + bx + c = 0). Use the Quadratic formula to solve for x. \[x = -b \pm \frac{ \sqrt{b^2 - 4ac} }{ 2a }\] \[x = -5 \pm \frac{ \sqrt{5^2 - 4(1)(-13)} }{ 2(1) }\]\[x = -5 \pm \frac{ \sqrt{25 + 52} }{ 2 }\]\[x = -5 \pm \frac{ \sqrt{77} }{ 2 } \]\[x = -5 \pm \frac{ 8.77}{ 2 }\]\[x = -5 \pm 4.385\] \[x = -5 + 4.835 =-0.615 \] \[x = -5 - 4.835 =-9.835 \] So: x = -0.615 and x = -9.835

OpenStudy (anonymous):

x^2-4x+5x-20=7 x^2+x-13=0 now find x.

OpenStudy (anonymous):

I made a mistake when subtracting 7. Okie so... x^2 + x - 20 = 7 Subtract 7. x^2 + 5x - 27 = 0 ^That is a quadratic equation (ax^2 + bx + c = 0). Use the Quadratic formula to solve for x. -------------------------------------- \[x = -b \pm \frac{ \sqrt{b^2 - 4ac} }{ 2a }\]\[x = -5 \pm \frac{\sqrt{ 5^2 - 4(1)(-27)} }{ 2(1) } \]\[x = -5 \pm \frac{\sqrt{ 25 + 108)} }{ 2 }\]\[x = -5 \pm \frac{ \sqrt{133} }{ 2 }\]\[x = -5 \pm \frac{ 11.53 }{ 2 }\]\[x = -5 \pm 5.765\] \[x = -5 + 5.765 = 0.765\] \[x = -5 - 5.765 = -10.765\] So: x = 0.765 and x = -10.765

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)